Tyrion:

HELP!!!! Describe how to transform the quantity of the third root of x to the fourth power, to the fifth power into an expression with a rational exponent. Make sure you respond with complete sentences.

Tyrion:

@snowflake0531

snowflake0531:

\[\sqrt[3]{x}\] so to make this into x, we do \[\sqrt[3]{x} ^3\] so if you want to change it to the fourth power, 4+3 equals 7, and \[\sqrt[3]{x} ^7 = x^4\] altho i'm not toosure of what hte question is asking...

Tyrion:

idek dats y i posted it to maybe get help